OpenStudy (anonymous):

5 divided by the difference of a number and 11 equals the quotient of 10 and the sum of the number and 1. Find the number

OpenStudy (anonymous):

\[\frac{5}{x-11}=\frac{10}{x+1} \]\[x=23 \]